Main Menu Contents Grades Course Contents » Week 5: Dynamics Weighing yourself in an Elevator Timer Notes A physicist (mass 95 kg) is standing on a scale in an elevator. While he is going down, the elevator is speeding up with 2.4 m/s2. What does the scale show? 874㎏ Submit Answer Tries 3/12 Previous Tries Post Discussion
Answer #1

Force = Mass*(Acceleration due to gravity)

Force = m*g

if the lift is accelerated upward then there will be a pseduo force of same magnitude as acceleration acting downward on the scale.

Force (F) = m*g

F = 95*(9.8+2.4)

F = 1159 N